### v1

# thanks: These authors contributed equallythanks: These authors contributed equally

Abstract
Fast, reliable logical operations are essential for the realization of useful quantum computers [1, 2, 3], as they are required to implement practical quantum algorithms at large scale. By redundantly encoding logical qubits into many physical qubits and using syndrome measurements to detect and subsequently correct errors, one can achieve very low logical error rates. However, for most practical quantum error correcting (QEC) codes such as the surface code, it is generally believed that due to syndrome extraction errors, multiple extraction rounds—on the order of the code distance $d$—are required for fault-tolerant computation [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Here, we show that contrary to this common belief, fault-tolerant logical operations can be performed with constant time overhead for a broad class of QEC codes, including the surface code with magic state inputs and feed-forward operations, to achieve “algorithmic fault tolerance”. Through the combination of transversal operations [7] and novel strategies for correlated decoding [15], despite only having access to partial syndrome information, we prove that the deviation from the ideal measurement result distribution can be made exponentially small in the code distance. We supplement this proof with circuit-level simulations in a range of relevant settings, demonstrating the fault tolerance and competitive performance of our approach. Our work sheds new light on the theory of quantum fault tolerance, potentially reducing the space-time cost of practical fault-tolerant quantum computation by orders of magnitude.
Quantum computers have the potential to solve certain computational problems much faster than their classical counterparts [1, 16]. Since most known applications require quantum computers with extremely low error rates, quantum error correction (QEC) and strategies for fault-tolerant quantum computing (FTQC) are necessary. These methods encode logical quantum information into a QEC code involving many physical qubits, such that the lowest weight logical error has weight equal to the code distance $d$ and is therefore unlikely.
Performing large-scale computation, however, comes with significant overhead [16, 2]. By performing syndrome extraction (SE), one can reveal error information and use a classical decoder to correct physical errors in software and interpret logical measurement results. However, in the presence of noisy syndrome measurements [4, 5, 6, 7, 10], one typically requires a number of SE rounds that scales linearly in $d$, i.e., $\Theta(d)$  [^1] The notation $g(x)=\Theta(f(x))$ indicates that two functions $f(x)$ and $g(x)$ have the same asymptotic scaling with $x$, or more precisely, that there exists some constants $c_{1}$ and $c_{2}$ such that $c_{1}f(x)\leq g(x)\leq c_{2}f(x)$ for sufficiently large $x$. (see Fig. 1(a)). This is the case, for example, for the celebrated surface code [8, 9, 10], one of the leading candidates for practical FTQC due to its simple 2D layout and competitive error thresholds. In typical compilations based on lattice surgery or braiding [11, 12, 13, 14, 18], each logical operation requires $\Theta(d)$ SE rounds, thus incurring a space-time volume per logical operation of $\Theta(d^{3})$. This reduces the logical clock speed by a factor proportional to the code distance, typically on the order of 10 –100 [16, 14]. The same considerations also apply when performing logical operations with many quantum low-density parity-check (QLDPC) codes [19, 20]. While there have been various efforts at addressing this challenge [5, 21], these alternative approaches introduce higher hardware complexity [22, 23, 20, 24] or necessitate certain properties of the underlying codes, such as the single shot QEC property [25, 26, 27, 28, 29], often incurring a trade-off between space and time when executing logical operations [30, 16, 2].
Figure 1: Algorithmic fault tolerance. (a) Conventional FT analysis separately examines each gadget (red boxes) in the circuit and ensures they are individually FT [4, 7, 31]. This requires $\Theta(d)$ syndrome extraction (SE) rounds to achieve FT. (b) Algorithmic FT directly uses all accessible syndrome information up to a logical measurement (blue box), and guarantees FT of the measurement result, even if the gadgets are not individually FT and if future syndrome information is not yet accessible (partial decoding). We realize algorithmic FT through transversal operations, and only require a single SE round per logical operation, thus allowing constant time implementations of logical operations.
We introduce and develop a novel approach to FTQC that we refer to as “algorithmic fault tolerance”, and show that it can lead to a substantial reduction in space-time cost. We focus on transversal implementations of Clifford circuits [7, 32] with magic state inputs and feed-forward [33], thereby allowing universal quantum computation. Such transversal gate capabilities have already been demonstrated in multiple hardware platforms, such as neutral atoms and trapped ions [34, 35, 36]. We show that contrary to the common belief, for any Calderbank-Shor-Steane (CSS) QLDPC code [6, 37], these operations can be performed fault-tolerantly with only constant time overhead per operation, provided that decoding can be implemented efficiently. The key idea is to consider the fault tolerance of the algorithm as a whole (Fig. 1(b)) [38, 39, 40]. We achieve this by performing correlated decoding [15, 36, 30] despite only having access to partial syndrome information, and ensuring consistency in the presence of magic states and feed-forward via additional operations in software. We verify such algorithmic fault tolerance through a combination of proofs and circuit-level numerical simulations of our protocol, including a simulation of state distillation factories [33, 13], finding very little change to physical error thresholds. Specializing to the surface code, our results reduce the per-operation time cost from $\Theta(d)$ to $\Theta(1)$, including for Clifford operations used in magic state distillation. Note that unlike methods that trade space for time, our techniques represent a direct reduction in space-time volume, which is usually the ultimate quantity of interest.
Algorithmic Fault Tolerance via Transversal Operations
We focus on transversal Clifford circuits with magic state inputs, where Clifford operations are implemented with a depth-one quantum circuit (Methods). This is interleaved with SE rounds using ancilla qubits, which reveal error information on the data qubits and enable error correction. In addition to transversal gates [7], we refer to preparation of data qubits in $|0\rangle$ followed by one SE round as transversal state preparation, and $Z$ basis measurement of all data qubits as transversal measurement. To achieve universality, we allow teleporting in low-noise magic states with feed-forward operations based on past measurement results, and use the same Clifford operations above to prepare high quality magic states via magic state distillation [33]. We make use of CSS QLDPC codes, where each data or ancilla qubit interacts with a constant number of other qubits, and each stabilizer generator consists of all $X$ or all $Z$ operators [37, 6]. Within this setting, our key result can be formulated as the following theorem:
Theorem 1 (informal): Exponential error suppression for constant time transversal Clifford operations with any CSS QLDPC code. For a transversal Clifford circuit with low-noise magic state inputs and feed-forward operations, that can be implemented with a given CSS QLDPC code family $\mathcal{Q}_{d}$ of growing code distance $d$, there exists a threshold $p_{\mathrm{th}}$, such that if the physical error rate $p<p_{\mathrm{th}}$ under the basic model of fault tolerance [5], then our protocol can perform constant time logical operations, with only a single SE round per operation, while suppressing the total logical error rate as $P_{L}=\exp(-\Theta(d_{n}))$.
The formal theorem statement and the corresponding proof can be found in Supplementary Materials [41]. Our analysis assumes the basic model of fault tolerance [5]. In particular, we consider the local stochastic noise model, where we apply depolarizing errors on each data qubit every SE round and measurement errors on each SE result, with a probability that decays exponentially in the weight of the error event. This can be readily generalized to circuit-level noise by noting the bounded error propagation for constant depth SE circuits in QLDPC codes. We also assume the most likely error (MLE) decoder and fast classical computation (Methods). Finally, we assume that all code patches are identical, and the number of qubit locations within a code patch that any given qubit can be coupled to via transversal gates is bounded by some constant $t$, in order to control error propagation.
A key observation is that by considering the algorithm as a whole and leveraging the deterministic propagation of errors through transversal Clifford circuits, one can use the surrounding syndrome history to correct for noisy measurements (Fig. 1(b)). This correlated decoding technique has been shown to enable $\Theta(1)$ SE rounds for Clifford circuits without feed-forward [15]. However, a key component of many schemes for achieving universality is magic state teleportation, which crucially relies on the ability to realize feed-forward operations.
As illustrated by the example shown in Fig. 2(a), such feed-forward operations require on-the-fly interpretation of logical measurements, followed by a subsequent conditional gate, when only a subset of the logical qubits have been measured. As we do not yet have future syndrome information on the unmeasured logical qubits, one may be concerned that this can lead to an incorrect assignment of logical measurement results. Indeed, prior work analyzing circuits with magic states assumed that at least $d$ SE rounds separated state initialization and measurements or out-going qubits [30, 42, 43]. As shown in Fig. 2(b) for the $\Theta(1)$ SE round case, with new syndrome information, one may end up concluding a different measurement result, which leads to an incorrect feed-forward operation.
Surprisingly, we find that these inconsistencies can be accounted for in classical processing, with a reinterpretation of subsequent measurement results (Fig. 2(c), Pauli frame updates). The inconsistent measurement result corresponds to an $\overline{X}$ operator applied right before the $\overline{Z}$ measurement. Tracing back, we can find an $\overline{X}$ operator on the $|{\overline{+}}\rangle$ initial state (Fig. 2(c)) which does not change the logical state but propagates through to apply $\overline{X}$ on the logical measurement, together with some other logical Pauli updates on the remaining logical qubits. These are stabilizers of the logical state, which leave the state invariant. Indeed, the fact that this measurement result can be affected by non-fault-tolerant state preparation implies that the measurement anti-commutes with the corresponding Pauli stabilizer, necessarily leading to a 50/50 random outcome that is not changed by a logical flip. Products of individual measurements can have nontrivial correlations only if they commute with all the Pauli stabilizers. Because they commute, however, they are also guaranteed to be insensitive to the state initialization errors.
Therefore, in the second step of our decoding procedure, we apply such Pauli operators on initial input states until the measurement results are consistent with the previous commitments (Fig. 2(c)). Beyond this specific circuit, the required pattern that leads to a consistent assignment can always be computed efficiently by solving a linear system of equations (Methods). In practice, subsets of measurements in which all measurement products are 50/50 random can be classically assigned in advance, with the future measurements determined through the above procedure to ensure consistency. This also implies that decoding of certain measurements can be delayed until joint products need to be determined, and some assignments can be performed deterministically in specific cases such as state distillation (Methods).
Our protocol that leads to Theorem 1 thus consists of two main steps: correlated decoding based on partial syndrome information, and application of logical stabilizers to guarantee consistency between multiple decoding rounds (Fig. 2).
Figure 2: Illustration of decoding strategy. (a) Logical quantum circuit with measurement and feed-forward. All logical operations are transversal and interleaved with a single SE round, instead of $d$ SE rounds. We must decode and commit mid-circuit to a measurement result for the bottom qubit, despite lacking complete syndrome information on the top two qubits (partial decoding). (b) With the measurement result of the bottom qubit, a feed-forward operation is applied, the remaining circuit is executed, and decoding is performed again on the whole circuit. The second decoding round may assign a different result to the bottom qubit, causing an inconsistency in feed-forward operations. (c) To guarantee consistency, we apply an $\overline{X}$ operator on the $|{\overline{+}}\rangle$ initial state of the middle qubit, which acts trivially on $|{\overline{+}}\rangle$, but changes the interpreted logical measurement result $M_{a}$ to be consistent with before. This also leads to a re-interpretation of the logical measurement result $M_{2}$.
We now sketch the intuition behind our proof of Theorem 1. There are two types of logical errors that may occur with our protocol. The first, a heralded inconsistency error, occurs when we are not able to find a set of operators to apply that yield the same outcome as previously committed measurement results. The second, a regular error, occurs when an erroneous logical operator is applied that results in a different measurement distribution.
Because imperfect readout during transversal measurements are equivalent to data qubit errors followed by perfect measurements, transversal measurements produce reliable syndrome information. Intuitively, this prevents individual errors from leading to high-weight corrections on the logical qubits we measure, the main reason for needing $d$ SE rounds in typical FT state initialization protocols. At the same time, the use of correlated decoding, together with the structured error propagation through transversal Clifford gates, allow us to propagate this syndrome information and correct relevant errors happening throughout the circuit. With these observations, we prove that for either type of logical error to occur, the total Pauli weight $s$ of physical error and subsequent correction in a connected cluster must satisfy $s=\Theta(d)$, which has probability $p^{s/2}$ under the MLE decoder. Finally, we count the number of such connected clusters of size $s$, which scales as $(v\mathfrak{e})^{s}$, where $\mathfrak{e}$ is the natural base and $v$ is a constant upper-bounding the error connectivity for a QLDPC code. The combined probability of an error thus scales as
$P_{\mathrm{err}}\propto p^{s/2}(2v\mathfrak{e})^{s}=\quantity(2v\mathfrak{e}\sqrt{p})^{\Theta(d)}\rightarrow 0$ (1)
when the physical error rate is sufficiently low $p\ll 1/(v\mathfrak{e})^{2}$ (the factor of 2 comes from a combinatorial sum), thereby establishing the existence of a threshold and exponential error suppression.
Specializing to the surface code and utilizing the full transversal Clifford gate set accessible to the surface code (Methods), an immediate corollary of our main theorem is a threshold result for performing constant time logical operations with an arbitrary transversal Clifford circuit. This result supports universal quantum computing when we allow magic state inputs prepared with sufficiently low noise.
Preparing high quality magic state inputs, in turn, can be performed simply with the same Clifford operations and easy-to-prepare non-fault-tolerant magic states [44, 45, 46], a procedure known as magic state distillation [33] (see ED Fig. 3). We expect that the same algorithmic FT approach described above achieves a $\Theta(d)$ speed-up in distillation time as well. The distillation factory and main computation can then be combined by applying our decoding approach to the joint system. In Methods and Supplementary Information, we further describe an extension of our results to the case of single-shot code patch growth, relevant to practical distillation factories [47, 48]. Taken together, these results provide a theoretical foundation for our factor of $\Theta(d)$ improvement in logical clock speed compared to standard FT approaches for universal quantum computation.
Figure 3: Numerical verification of fault tolerance. (a) Simulation of circuit with repeated $\overline{ZZ}$ measurement (inset), where we commit mid-circuit to each measurement result of the logical ancilla using only the syndrome information up to that point. The total logical error rate as a function of circuit-level physical error rate $p$, for varying code distance $d$, shows clear threshold behavior. (b) Heralded error rate with and without the second step of our decoding strategy, as a function of code distance and for different physical error rates, for the same circuit as (a). Only with both steps do we observe exponential suppression of the logical error rate. (c) Comparison of two different methods for logical state preparation between three rotated surface codes and subsequent teleportation, for fixed circuit noise $p=0.3\%$. We use transversal gates (left) and lattice surgery (right), in both cases with only a single SE round. (d) With transversal gates, the error rate decreases exponentially with the code distance. With a single round of lattice surgery, the error rate instead increases linearly with code distance, as a single stabilizer measurement error affects the logical $\overline{ZZ}$ measurement result.
Competitive Numerical Performance
Figure 4: $|{\overline{Y}}\rangle$ state distillation factory. (a) Illustration of a $|{\overline{Y}}\rangle$ state distillation factory based on the [[7,1,3]] Steane code, consisting of state initialization, layers of transversal CNOTs, followed by a teleported $\overline{S}$ gate. Each operation involves only a single SE round. Two of the CNOTs in the first layer act trivially and can be omitted. (b) The $|{\overline{Y}}\rangle$ resource state is prepared via state injection at the first level, and via the first-level factory for the second level. (c) 1-level factory output state infidelity as a function of input state infidelity, for fixed circuit noise $p=0.1\%$ and varying levels of artificially injected $Z$ errors. The ideal curve is calculated assuming the gate operations in the factory are perfect. (d) Performance for one and two rounds of distillation, showing good agreement with the expected scaling.
We now turn to circuit-level simulations of our protocol to numerically evaluate its performance [39], and contrast it with existing methods. We consider various test cases of our approach that also serve as key subroutines in large-scale algorithms.
We first consider a simple circuit with intermediate logical measurements (inset of Fig. 3(a)). In this example, two logical qubits are transversally initialized in $|{\overline{+}}\rangle$, and an ancilla logical qubit is used to measure the $\overline{Z}\overline{Z}$ correlation a total of eight times, before the two logical qubits are transversally measured in the $\overline{Z}$ basis. While individual logical measurement results are random, a correct realization of this circuit should yield the same result for $\overline{Z}\overline{Z}$ each time, which in turn should be consistent with the final logical measurement results. We employ our algorithmic FT protocol to decode the circuit up to each logical measurement using only the syndrome information accessible at that point. We use the rotated surface code, a circuit-level depolarizing noise model [49, 15], a MLE decoder based on integer programming [50, 15], and employ the two-step process described above (see Supplementary Information).
Figure 3(a-b) show the results of numerical simulations. We find that the total logical error rate, defined as the probability that a logical error of either type mentioned above happened anywhere in the circuit, shows characteristic threshold behavior, with an estimated threshold $\gtrsim 0.85\%$. As an SE round involves four layers of CNOT gates, while the transversal CNOT only involves a single layer, the effective error rate is dominated by SE operations, hence it may be expected that the threshold is close to the circuit-noise memory threshold. The number of SE rounds can be further optimized: for example, in Ref. [15], performing one SE round every four gate layers minimized the space-time cost per CNOT, suggesting that the practical improvement may be $\gtrsim 2d$ in some regimes  [^2] Four transversal gates and one SE round require 8 CNOT gates in total, compared to $416d+$ 4 with $d$ SE rounds.. In Fig. 3(b), we further compare the scaling of heralded failure rates in the presence and absence of the second step of our decoding procedure, as a function of code distance $d$. We find that this additional step is crucial to achieve exponential suppression with the code distance.
We now contrast our approach with lattice surgery in a similar setting [11, 12, 52, 18]. We consider the logical circuit in Fig. 3(c), where a GHZ state preparation circuit is followed by teleportation of the GHZ state to another set of logical qubits, and then measurement in the $\overline{Z}$ basis [41]. Using transversal gates with only a single SE round during $|{\overline{+}}\rangle$ and $|{\overline{0}}\rangle$ state preparation, and decoding each logical measurement with only accessible information at that stage, we find that the logical error rate decreases exponentially with the code distance, consistent with our FT analysis. In contrast, state preparation based on a single round of lattice surgery [52], which involves performing syndrome extraction with a larger code patch and then splitting it into three individual logical qubits, does not yield improved logical error rate as the code distance increases, as a single error can lead to incorrect inference of the $\overline{Z}\overline{Z}$ correlation of the GHZ state (Supplementary Information). Unlike transversal measurements, logical information here is contained in noisy stabilizer products, which require repetition to reliably infer.
Next, we simulate a state distillation factory. In order to perform a classical simulation of a full factory, we focus on distillation of the $|{\overline{Y}}\rangle=\overline{S}|{\overline{+}}\rangle$ state (Fig. 4(a)), which allows the easy implementation of $\overline{S}$ gates in the surface code. Since this circuit has a similar structure to the practically relevant $|{\overline{T}}\rangle$ magic state distillation factories (Methods, ED Fig. 3), we expect them to have similar performance. We fix the error rate of the circuit to $p=0.1\%$, and vary the input infidelity $P_{\mathrm{in}}$ in Fig. 4(c). Examining the output $|{\overline{Y}}\rangle$ of a one-level factory, we find that as the code distance is increased, the output logical error rate $P_{\mathrm{out}}$ approaches the fidelity expected for ideal Clifford logical gates in the factory $P_{\mathrm{out}}=7P_{\mathrm{in}}^{3}+O(P_{\mathrm{in}}^{4})$ (see Methods for the full expression), across the explored fidelity regime.
Finally, we simulate the logical error rate for a two-level $|{\overline{Y}}\rangle$ state distillation factory, involving a total of 113 logical qubits, where the output $|{\overline{Y}}\rangle$ states of a $d_{1}=5$ factory is fed into a second factory with $d_{2}=9$, with the distance chosen such that the logical error is dominated by the input state infidelity. As shown in Fig. 4(d), the logical error rates at each level of the distillation procedure are consistent with that expected based on the ideal factory formula (Methods), confirming that our approach is FT. Since the state injection procedure is agnostic to the particular state that is injected, we expect that our results will readily generalize to the setting of $|{\overline{T}}\rangle$ magic state factories.
Discussion and Outlook
Transversal operations and correlated decoding were recently found to be highly effective in experiments with reconfigurable neutral atom arrays [36]. The principles of algorithmic fault-tolerance described here are the core underlying mechanisms of these observations, such as correlated decoding of a logical Bell state [36], and our results here indicate that the same techniques allow for $\Theta(d)$ time reduction for universal computation. While recent work has provided strong evidence that this reduction might be possible for circuits consisting purely of Clifford gates and Pauli basis inputs [15], up to now it has generally been believed that this conclusion does not hold when performing universal quantum computation [30, 42, 43], which crucially relies on the use of magic states and feed-forward operations. The present work not only demonstrates that this $\Theta(d)$ time cost reduction is broadly applicable to universal quantum computing, but also provides a theoretical foundation for it through mathematical fault tolerance proofs.
Although our analysis focused on the use of an MLE decoder, our numerical simulations suggest that algorithms with polynomial runtime can still achieve a competitive threshold [41], and the development of improved, parallel correlated decoders is an important area of future research (Methods). Taking into account the decoding time overhead, we may eventually need to insert more SE rounds to simplify decoding or wait for decoding completion [53], as is also needed for FT protocols that rely on single-shot quantum error correction [25]. In that case, we still expect a significant practical saving over existing schemes. In light of recent experimental advances  [36], a full compilation and evaluation of the space-time savings in parallel reconfigurable architectures such as neutral atom arrays is an important next step. Finally, it will be interesting to investigate how these results can be combined with recent progress toward constant-space-overhead quantum computation [5, 54, 55, 56, 23, 20, 27, 29] or generalized to transversal non-Clifford gates [57, 30, 58, 59, 60, 61, 62], in order to further reduce the space-time volume of large-scale quantum computation.
Author Contributions
H.Z. formulated the decoding strategy and developed an initial proof sketch through discussions with C.Z., M.C., D.B., C.D., S.-T.W., A.K., and M.D.L.. C.Z., M.C., H.Z., and H.H. performed numerical simulations. H.Z., A.K., C.Z., M.C., and C.D. proved the fault tolerance of the scheme. All authors contributed to writing the manuscript.
Acknowledgements
We acknowledge helpful discussions with G. Baranes, P. Bonilla, E. Campbell, S. Evered, S. Geim, L. Jiang, M. Kalinowski, A. Krishna, S. Li, D. Litinski, T. Manovitz, N. Maskara, Y. Wu, and Q. Xu. We would particularly like to thank C. Pattison for early discussions and suggesting the simulation of the $|{\overline{Y}}\rangle$ state distillation factory, and J. Haah for stimulating discussions and deep insights. We acknowledge financial support from IARPA and the Army Research Office, under the Entangled Logical Qubits program (Cooperative Agreement Number W911NF-23-2-0219), the DARPA ONISQ program (grant number W911NF2010021), the DARPA IMPAQT program (grant number HR0011-23-3-0012), the Center for Ultracold Atoms (a NSF Physics Frontiers Center, PHY-1734011), the National Science Foundation (grant number PHY-2012023 and grant number CCF-2313084), the Army Research Office MURI (grant number W911NF-20-1-0082), DOE/LBNL (grant number DE-AC02-05CH11231). M.C. acknowledges support from Department of Energy Computational Science Graduate Fellowship under award number DE-SC0020347. D.B. acknowledges support from the NSF Graduate Research Fellowship Program (grant DGE1745303) and The Fannie and John Hertz Foundation. This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA). The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.